# SocialTrack: Multi-Object Tracking in Complex Urban Traffic Scenes Inspired by Social Behavior

### v2

@@ Abstract
Index Terms:
Multi-object tracking, spatio-temporal memory fusion, multi-scale context aggregation, group motion compensation, UAV.
I Introduction
As one of the core tasks in the field of machine vision, multi-object tracking (MOT)[1] holds significant importance in the context of urban intelligent transportation systems. The main objective is to maintain continuous motion trajectories for each detected object in a video sequence[2]. In urban traffic management, MOT technology not only requires efficient object detection but also necessitates powerful object association capabilities to address challenges such as occlusions, object crossings, and rapid motion in complex environments. With the rapid development of computer vision, hardware capabilities, and sensor technologies, MOT has shown immense potential in applications such as autonomous driving[3][4], urban traffic monitoring[5], intelligent transportation systems[6], and urban planning.
In urban traffic scenarios, the application of MOT spans a wide range, including the detection and tracking of vehicles, pedestrians, roads, and other elements. For example, in intelligent transportation systems, MOT technology can help with dynamic traffic flow monitoring, road congestion analysis, and traffic trend prediction[7]. With the rapid development of unmanned aerial vehicle (UAV) technology[8], UAVs, with their high maneuverability, convenient deployment, and flexible viewpoints, have become important tools for urban traffic monitoring[9][10]. UAVs equipped with cameras can be used for monitoring urban traffic conditions[11], detecting traffic accidents, and assessing road safety[12]. However, the complex urban environment, especially highly dynamic traffic scenarios, presents new challenges for object tracking[13], such as target scale variation, occlusions, nonlinear crossing motions, and motion blur.
@@ I Introduction
To address the issue of trajectory instability, recent research has primarily focused on appearance learning and association design[16]. In terms of appearance learning, MeMOT[17] retains a large spatiotemporal memory to store the identity embeddings of the tracked targets, adaptively referencing and aggregating useful information from memory as needed. The work in[18] explores diversified fine-grained representations of target appearance using high feature resolution and precise semantic information. Although these methods enhance the discriminability of appearance features, they are ineffective in handling small targets with missing appearance features and require additional computational resources, which hinders real-time tracking. In terms of association design, the work in[19] combines appearance features with a simple motion model to achieve strong tracking results. SparseTrack[20] proposes a pseudo-depth estimation method for deep cascading matching to improve occlusion handling in target association. BoT-SORT[21] employs dedicated camera motion compensation (CMC) technology to achieve frame alignment and handle dynamic scenes. Similarly, OC-SORT[22] constructs virtual trajectory segments during occlusions based on target observations to fill in missing trajectory information. Although these association strategies are effective for multi-object tracking, they are often designed for specific datasets or require fine-tuning of hyperparameters for particular scenarios.
@@
Unlike existing methods, this paper proposes a framework for multi-object tracking of small targets in complex urban traffic environments—SocialTrack. This framework integrates several strategies, including detector improvements, trajectory filtering, group prior modeling, and spatiotemporal prediction, to systematically enhance the tracking performance of small targets under challenging conditions such as occlusions, dense scenes, and motion blur. First, we introduce a multi-scale feature enhancement mechanism and a context encoding-decoding perception module into the detector architecture, called SOFEPNet, which effectively improves the response strength of small targets in the feature map. Next, to improve the accuracy of trajectory state estimation, we propose the Velocity Adaptive Cubature Kalman Filter (VACKF). This method incorporates a velocity dynamic modeling mechanism into the traditional filter framework and uses a nonlinear volumetric representation of the state space to improve the modeling of fast-moving small targets and nonlinear trajectory variations. To further enhance performance in dense object environments, we introduce a Group Motion Compensation Strategy (GMCS). By constructing a motion consistency modeling network in the local neighborhood, it provides stable motion references for low-quality trajectories, enabling state compensation and correction. Finally, we propose a Spatio-Temporal Memory Prediction (STMP) module, which is specifically designed to infer the future state of low-quality trajectories by leveraging the target’s own historical trajectory information in the absence of external observations or when the group compensation fails. Experimental results on the UAVDT and MOT17 public datasets demonstrate that the proposed SocialTrack outperforms existing state-of-the-art methods in terms of performance.
@@
The rest of the paper is organized as follows: Section II discusses related work on tracking frameworks; Section III presents the overall architecture of the SocialTrack and the proposed improvement strategies; Section IV provides comprehensive experimental validation of the proposed improvements; and Section V concludes the paper.
II Related Work
II-A MOT
@@
Single-stage. Single-stage object trackers perform both detection and motion/appearance feature extraction simultaneously, ultimately outputting the target’s position and unique identity. CenterTrack[32] focuses on the center of the target, directly predicting the target’s position and motion offset using a network. TraDeS[33] uses tracking clues to end-to-end assist detection, constructing a cost matrix to infer motion offsets for target tracking and propagating previous target features based on offsets to improve detection performance. In terms of appearance feature information, Swin-JDE[34] employs neural network learning for upsampling, enhancing the spatial information of the feature map so that a single network can simultaneously perform detection and re-identification tasks. FairMOT[35] introduces a re-identification branch to predict target features and jointly optimizes the detection and re-identification tasks through a carefully designed network structure. Chained-Tracker[36] integrates object detection, feature extraction, and data association into a single end-to-end solution for the first time. HomView-MOT[37] incorporates homographic matching and viewpoint-centric concepts, achieving cross-view target association through multi-view homogeneity learning. U2MOT[38] improves feature consistency by leveraging uncertainty, proposing an unsupervised multi-object tracking framework based on uncertainty metrics and trajectory segment-guided enhancement strategies. UCMCTrack[39] effectively handles challenges caused by camera motion by applying unified camera motion compensation parameters throughout the video sequence and introducing position mapping. TrackFormer[40] uses attention mechanisms to autoregressively track target trajectories in video sequences, without requiring additional graph optimization or appearance modeling. MOTR[41] extends DETR[42] by introducing ”tracking queries” and combines small trajectory-aware label assignment with a temporal aggregation network to effectively model temporal relationships and improve tracking accuracy. TrackingMamba[43] adopts a single-stream architecture with a Vision Mambacit[44] backbone, combining global feature extraction and long-range dependency modeling, significantly reducing computational overhead.
II-B Data association
Data association is a critical step in Multi-Object Tracking (MOT). It works by evaluating the similarity between existing trajectories and the detection results of the current frame, thereby matching targets across different frames and assigning a unique ID to each target[45]. MOTDT[46] collects candidate targets from both detection and tracking outputs and utilizes a scoring function based on fully convolutional neural networks to select the optimal target for matching in real-time. QDTrack[47] proposes a quasi-dense similarity learning method by densely sampling hundreds of region proposals on a pair of images for contrastive learning, and performs simple nearest-neighbor search using the obtained feature space. SparseTrack[20] improves data association performance in crowded and occluded scenarios by using pseudo-depth estimation and the Deep Cascade Matching (DCM) strategy, which transforms dense target sets into sparse target subsets. Unlike the aforementioned methods, we leverage group motion pattern learning to infer the reliable motion states of all targets and apply this knowledge to existing trajectories.
III Methodology
III-A Overview of the Framework
@@ 31     $T_{low}^{center}=Cascade\_LSTM({T_{low}})$ ${T_{low}}\leftarrow VACKF\_update(T_{low}^{center})$
Unlike existing methods such as[23][25][26] that focus solely on high-quality trajectories, we allocate more resources to handling low-quality tracking trajectories, as these are the main causes of identity switches and trajectory drift. For unmatched low-quality trajectories, we design a Group Motion Compensation Strategy (GMCS) as prior knowledge to update the trajectory states, preventing error accumulation. When the motion compensation strategy fails, a Spatio-Temporal Memory Prediction (STMP) module is employed to predict the object’s position at the next time step using historical trajectory information. The final output of the algorithm is a set of trajectories $T$, where each trajectory includes cross-frame bounding box coordinates and the unique ID of the object. The overall architecture of SocialTrack is demonstrated in Fig. 1.
III-B Small Object Feature Efficient Perception Network
UAVs typically operate at high altitudes, with onboard cameras providing wide fields of view. However, the targets captured in the images are often small, with low resolution, and lack shape, structure, and texture information. Moreover, UAVs encounter various shooting environments, such as changes in viewpoint (e.g., top-down, side-view), lighting conditions (e.g., sunlight, night), and weather conditions (e.g., fog, rain). Conventional object detection methods utilize backbone networks like ResNet[48] to extract features at different scales, then perform simple multi-scale feature fusion using feature pyramid networks, ultimately producing detection results at different scales to tackle multi-scale object detection tasks. However, these networks face several issues when handling small targets: 1) excessive downsampling of high-level features results in a significant loss of crucial information for small targets; 2) low-level features lack semantic information, leading to inadequate representation of small target features; 3) inefficient feature fusion mechanisms prevent the effective enhancement of small target features.
@@ III-B Small Object Feature Efficient Perception Network
Feature upsampling is a critical step in restoring feature resolution in dense prediction models. Specifically, DED mainly consists of a dynamic upsampling splicing structure (DUSS), as shown on the right side of Fig. 2. DUSS is composed of a 3×3 convolution, DySample, and the C3 module. Given a feature, after passing through a regular convolution, it is input into DySample for dynamic upsampling of feature maps, resulting in an upsampled feature of size:
@@
III-C Velocity Adaptive Cubature Kalman Filter
@@
III-D Group Motion Compensation Strategy
@@
III-E Spatio Temporal Memory Prediction
@@
IV Experimental Results and Analysis
@@ IV-A Datasets
UAVDT: UAVDT contains 50 video sequences for multi-object tracking tasks, with 30 sequences for training and 20 sequences for validation. The UAVDT dataset has the following characteristics: 1) High density: wide viewing angles with a large number of targets; 2) Small targets: high shooting altitude with small target sizes; 3) Camera motion: the UAV flies at high speed with camera rotation. The video sequences are stored in JPG format with resolutions of 1080×540 pixels or 960×540 pixels and are captured in various complex scenarios, such as plazas, toll stations, highways, and intersections. UAVDT includes three types of natural environments: daytime, nighttime, and foggy conditions, with complex camera viewpoint variations, such as front view, top-down view, and bird’s-eye view, which present a significant challenge to multi-object tracking algorithms. UAVDT includes three object categories: cars, trucks, and buses, with a severe class imbalance: cars make up 90% of all instances, while trucks and buses each account for less than 5%.
MOT17: MOT17 is a classic and challenging dataset for multi-object tracking, focusing on pedestrian multi-object tracking. It contains 14 video sequences, with 7 for training and 7 for testing, at a resolution of 1920×1080. The main challenges of MOT17 are dense scenes, occlusion, and appearance similarity interference. The image sequences feature dense pedestrians, with severe occlusion and high appearance similarity between objects. We split the training sequences evenly, using the first half for training and the latter half for validation, focusing on the pedestrian category in the multi-object tracking task.
IV-B UAVDT Correction
@@ IV-D Implementation Details
TABLE I: Comparison of SocialTrack with Other Methods on UAVDT Dataset. The Last Row Shows Our Proposed SocialTrack, Which Obtained SOTA Performance in Almost All Categories.
@@
IV-E Comparison with Other State-of-the-Art Methods
To comprehensively evaluate the overall performance of the proposed SocialTrack algorithm in multi-object tracking tasks, we conducted a systematic comparison with current mainstream multi-object tracking methods on the challenging UAVDT test set. The comparison included traditional trackers (such as SORT, DeepSORT), end-to-end methods (such as JDE, FairMOT), and recently proposed advanced algorithms (such as OC-SORT, ByteTrack, BoT-SORT). The comparison results are shown in Table I. From the results, it can be seen that SocialTrack significantly outperforms existing methods on several key evaluation metrics, achieving the current best performance.
@@ IV-E Comparison with Other State-of-the-Art Methods
TABLE II: Complete Results of the Ablation Study on UAVDT.

### v3

@@ Abstract
keywords:
Multi-object tracking, spatio-temporal memory fusion, multi-scale context aggregation, group motion compensation, UAV
1 Introduction
As one of the core tasks in the field of machine vision, multi-object tracking (MOT) [52] holds significant importance in the context of urban intelligent transportation systems. The main objective is to maintain continuous motion trajectories for each detected object in a video sequence [42]. In urban traffic management, MOT technology not only requires efficient object detection but also necessitates powerful object association capabilities to address challenges [18] such as occlusions, object crossings, and rapid motion in complex environments. With the rapid development of computer vision, hardware capabilities, and sensor technologies, MOT has shown immense potential in applications such as autonomous driving [9, 53], urban traffic monitoring [27], intelligent transportation systems [43], and urban planning.
In urban traffic scenarios, the application of MOT spans a wide range, including the detection and tracking of vehicles, pedestrians, roads, and other elements. For example, in intelligent transportation systems, MOT technology can help with dynamic traffic flow monitoring, road congestion analysis, and traffic trend prediction [17]. With the rapid development of unmanned aerial vehicle (UAV) technology [40], UAVs, with their high maneuverability, convenient deployment, and flexible viewpoints, have become important tools for urban traffic monitoring [67, 35]. UAVs equipped with cameras can be used for monitoring urban traffic conditions [41], detecting traffic accidents, and assessing road safety [30]. However, the complex urban environment, especially highly dynamic traffic scenarios, presents new challenges for object tracking [20], such as target scale variation, occlusions, nonlinear crossing motions, and motion blur.
@@ 1 Introduction
To address the issue of trajectory instability, recent research has primarily focused on appearance learning and association design [22]. In terms of appearance learning, MeMOT [10] retains a large spatiotemporal memory to store the identity embeddings of the tracked targets, adaptively referencing and aggregating useful information from memory as needed. The work in [46] explores diversified fine-grained representations of target appearance using high feature resolution and precise semantic information. Although these methods enhance the discriminability of appearance features, they are ineffective in handling small targets with missing appearance features and require additional computational resources, which hinders real-time tracking. In terms of association design, the work in [49] combines appearance features with a simple motion model to achieve strong tracking results. SparseTrack [36] proposes a pseudo-depth estimation method for deep cascading matching to improve occlusion handling in target association. BoT-SORT [1] employs dedicated camera motion compensation (CMC) technology [26] to achieve frame alignment and handle dynamic scenes. Similarly, OC-SORT [11] constructs virtual trajectory segments during occlusions based on target observations to fill in missing trajectory information. Although these association strategies are effective for multi-object tracking, they are often designed for specific datasets or require fine-tuning of hyperparameters for particular scenarios.
@@
Unlike existing methods, this paper proposes a framework for multi-object tracking of small targets in complex urban traffic environments-SocialTrack. This framework integrates several strategies, including detector improvements, trajectory filtering, group prior modeling, and spatiotemporal prediction, to systematically enhance the tracking performance of small targets under challenging conditions such as occlusions, dense scenes, and motion blur. First, we introduce a multi-scale feature enhancement mechanism and a context encoding-decoding perception module into the detector architecture, called SOFEPNet, which effectively improves the response strength of small targets in the feature map. Next, to improve the accuracy of trajectory state estimation, we propose the Velocity Adaptive Cubature Kalman Filter (VACKF). This method incorporates a velocity dynamic modeling mechanism into the traditional filter framework and uses a nonlinear volumetric representation of the state space to improve the modeling of fast-moving small targets and nonlinear trajectory variations. To further enhance performance in dense object environments, we introduce a Group Motion Compensation Strategy (GMCS). By constructing a motion consistency modeling network in the local neighborhood, it provides stable motion references for low-quality trajectories, enabling state compensation and correction. Finally, we propose a Spatio-Temporal Memory Prediction (STMP) module, which is specifically designed to infer the future state of low-quality trajectories by leveraging the target’s own historical trajectory information in the absence of external observations or when the group compensation fails. Experimental results on the UAVDT and MOT17 public datasets demonstrate that the proposed SocialTrack outperforms existing state-of-the-art methods in terms of performance.
@@
The rest of the paper is organized as follows: Section 2 discusses related work on tracking frameworks; Section 3 presents the overall architecture of the SocialTrack and the proposed improvement strategies; Section 4 provides comprehensive experimental validation of the proposed improvements; and Section 5 concludes the paper.
2 Related Work
2.1 MOT
@@
Single-stage. Single-stage object trackers perform both detection and motion/appearance feature extraction simultaneously, ultimately outputting the target’s position and unique identity. CenterTrack [66] focuses on the center of the target, directly predicting the target’s position and motion offset using a network. TraDeS [56] uses tracking clues to end-to-end assist detection, constructing a cost matrix to infer motion offsets for target tracking and propagating previous target features based on offsets to improve detection performance. In terms of appearance feature information, Swin-JDE [51] employs neural network learning for upsampling, enhancing the spatial information of the feature map so that a single network can simultaneously perform detection and re-identification tasks. FairMOT [63] introduces a re-identification branch to predict target features and jointly optimizes the detection and re-identification tasks through a carefully designed network structure. Chained-Tracker [45] integrates object detection, feature extraction, and data association into a single end-to-end solution for the first time. HomView-MOT [25] incorporates homographic matching and viewpoint-centric concepts, achieving cross-view target association through multi-view homogeneity learning. U2MOT [32] improves feature consistency by leveraging uncertainty, proposing an unsupervised multi-object tracking framework based on uncertainty metrics and trajectory segment-guided enhancement strategies. UCMCTrack [59] effectively handles challenges caused by camera motion by applying unified camera motion compensation parameters throughout the video sequence and introducing position mapping. TrackFormer [38] uses attention mechanisms to autoregressively track target trajectories in video sequences, without requiring additional graph optimization or appearance modeling. MOTR [62] extends DETR [12] by introducing "tracking queries" and combines small trajectory-aware label assignment with a temporal aggregation network to effectively model temporal relationships and improve tracking accuracy. TrackingMamba [54] adopts a single-stream architecture with a Vision Mambacit [60] backbone, combining global feature extraction and long-range dependency modeling, significantly reducing computational overhead.
@@ 31     $T_{low}^{center}=Cascade\_LSTM({T_{low}})$ ${T_{low}}\leftarrow VACKF\_update(T_{low}^{center})$
2.2 Data association
Data association is a critical step in Multi-Object Tracking (MOT). It works by evaluating the similarity between existing trajectories and the detection results of the current frame, thereby matching targets across different frames and assigning a unique ID to each target [16]. [7] propose a linear projection method based on graph convolutional networks for effective correlation and classification of data. MOTDT [13] collects candidate targets from both detection and tracking outputs and utilizes a scoring function based on fully convolutional neural networks to select the optimal target for matching in real-time. QDTrack [44] proposes a quasi-dense similarity learning method by densely sampling hundreds of region proposals on a pair of images for contrastive learning, and performs simple nearest-neighbor search using the obtained feature space. SparseTrack [36] improves data association performance in crowded and occluded scenarios by using pseudo-depth estimation and the Deep Cascade Matching (DCM) strategy, which transforms dense target sets into sparse target subsets. Unlike the aforementioned methods, we leverage group motion pattern learning to infer the reliable motion states of all targets and apply this knowledge to existing trajectories.
3 Methodology
3.1 Overview of the Framework
Unlike existing methods such as [6, 55, 64] that focus solely on high-quality trajectories, we allocate more resources to handling low-quality tracking trajectories, as these are the main causes of identity switches and trajectory drift. For unmatched low-quality trajectories, we design a Group Motion Compensation Strategy (GMCS) as prior knowledge to update the trajectory states, preventing error accumulation. When the motion compensation strategy fails, a Spatio-Temporal Memory Prediction (STMP) module is employed to predict the object’s position at the next time step using historical trajectory information. The final output of the algorithm is a set of trajectories $T$, where each trajectory includes cross-frame bounding box coordinates and the unique ID of the object. The overall architecture of SocialTrack is demonstrated in Fig. 1.
3.2 Small Object Feature Efficient Perception Network
UAVs typically operate at high altitudes, with onboard cameras providing wide fields of view. However, the targets captured in the images are often small, with low resolution, and lack shape, structure, and texture information [24]. Moreover, UAVs encounter various shooting environments, such as changes in viewpoint (e.g., top-down, side-view), lighting conditions (e.g., sunlight, night), and weather conditions (e.g., fog, rain). Conventional object detection methods utilize backbone networks like ResNet [23] to extract features at different scales, then perform simple multi-scale feature fusion using feature pyramid networks, ultimately producing detection results at different scales to tackle multi-scale object detection tasks. However, these networks face several issues when handling small targets: 1) excessive downsampling of high-level features results in a significant loss of crucial information for small targets; 2) low-level features lack semantic information, leading to inadequate representation of small target features; 3) inefficient feature fusion mechanisms prevent the effective enhancement of small target features.
@@ 3.2 Small Object Feature Efficient Perception Network
Feature upsampling is a critical step in restoring feature resolution in dense prediction models. Specifically, DED mainly consists of a dynamic upsampling splicing structure (DUSS), as shown on the right side of Fig. 2. DUSS is composed of a $3\times 3$ convolution, DySample, and the C3 module. Given a feature, after passing through a regular convolution, it is input into DySample for dynamic upsampling of feature maps, resulting in an upsampled feature of size:
@@
3.3 Velocity Adaptive Cubature Kalman Filter
@@
3.4 Group Motion Compensation Strategy
@@
3.5 Spatio Temporal Memory Prediction
@@
4 Experimental Results and Analysis
@@ 4.1 Datasets
UAVDT: UAVDT contains 50 video sequences for multi-object tracking tasks, with 30 sequences for training and 20 sequences for validation. The UAVDT dataset has the following characteristics: 1) High density: wide viewing angles with a large number of targets; 2) Small targets: high shooting altitude with small target sizes; 3) Camera motion: the UAV flies at high speed with camera rotation. The video sequences are stored in JPG format with resolutions of $1080\times 540$ pixels or $960\times 540$ pixels and are captured in various complex scenarios, such as plazas, toll stations, highways, and intersections. UAVDT includes three types of natural environments: daytime, nighttime, and foggy conditions, with complex camera viewpoint variations, such as front view, top-down view, and bird’s-eye view, which present a significant challenge to multi-object tracking algorithms. UAVDT includes three object categories: cars, trucks, and buses, with a severe class imbalance: cars make up 90% of all instances, while trucks and buses each account for less than 5%.
MOT17: MOT17 is a classic and challenging dataset for multi-object tracking, focusing on pedestrian multi-object tracking. It contains 14 video sequences, with 7 for training and 7 for testing, at a resolution of $1920\times 1080$. The main challenges of MOT17 are dense scenes, occlusion, and appearance similarity interference. The image sequences feature dense pedestrians, with severe occlusion and high appearance similarity between objects. We split the training sequences evenly, using the first half for training and the latter half for validation, focusing on the pedestrian category in the multi-object tracking task.
4.2 UAVDT Correction
@@ 4.4 Implementation Details
4.5 Comparison with Other State-of-the-Art Methods
To comprehensively evaluate the overall performance of the proposed SocialTrack algorithm in multi-object tracking tasks, we conducted a systematic comparison with current mainstream multi-object tracking methods on the challenging UAVDT test set. The comparison included traditional trackers (such as SORT, DeepSORT), end-to-end methods (such as JDE, FairMOT), and recently proposed advanced algorithms (such as OC-SORT, ByteTrack, BoT-SORT). The comparison results are shown in Table 1. From the results, it can be seen that SocialTrack significantly outperforms existing methods on several key evaluation metrics, achieving the current best performance.
Table 1: Comparison of SocialTrack with other methods on the UAVDT dataset. The last row shows our proposed SocialTrack, which obtained SOTA performance in almost all categories.
@@ 4.5 Comparison with Other State-of-the-Art Methods
SMOT [14] 33.9 72.2 45 524 367 57112 166528 1752
IOUT [8] 36.6 72.1 32.7 534 357 42245 163881 993
CMOT [4] 36.9 74.7 57.5 664 351 69109 144760 1111
JDE [51] 39.5 73.5 55.3 624 442 - - 3124
FairMOT [63] 44.9 72.7 66.5 665 326 - - 2279
MDP [57] 43 73.5 61.5 647 324 46151 147735 541
SuperMOT [47] 50.6 75.7 70.6 674 227 60587 107538 407
MOTDT [13] 36.4 - 56.4 464 402 41505 196014 590
SORT [6] 39 74.3 43.7 484 400 33037 172628 2350
DeepSORT [55] 40.7 73.2 52.2 595 358 44868 155290 2061
OC-SORT [11] 52.1 - 69.8 535 192 37908 141340 286
BoT-SORT [1] 52.5 - 69.7 646 201 56067 121304 441
ByteTrack [64] 53 - 70.3 668 193 56021 119674 215
SFTrack [50] 55.3 - 71.3 682 184 49946 117224 284
SocialTrack (ours) 64.6 76.3 76.1 702 170 31133 89146 276
@@
4.6 Ablation Experiments
UAVDT: In this study, ByteTrack [64] was used as the baseline, and the effectiveness of each proposed module was validated through systematic ablation experiments. Table 2 shows the improvements in overall tracking performance due to four key modifications. First, to enhance the detector’s ability to perceive small targets, we designed a specialized detector for efficient small-target feature perception. The experimental results indicate that this module improved MOTA from 56.1 to 62.5, MOTP from 75.2 to 76.0, and IDF1 from 68.6 to 74.1, with significant progress in all three key metrics. This demonstrates that the improvement in detection quality plays a crucial role in the overall tracking performance. Additionally, IDSW decreased by 128, indicating that more accurate target detection also helps maintain identity consistency in the subsequent association process. Building on the detection results, we introduced the VACKF, which combines the advantages of the velocity-adaptive model and cubature Kalman filter for nonlinear state modeling and prediction. This module increased MOTA to 63.2, MOTP slightly to 76.1, and IDF1 to 74.7, while further decreasing IDSW to 304. This result demonstrates that the adaptive filtering mechanism significantly reduces trajectory estimation errors and identity switching frequency in scenes with frequent occlusions or complex motion, thereby enhancing the robustness and continuity of the tracking system. Next, we introduced the GMCS, which leverages the motion prior information from neighboring trajectories in the group to provide robust relative state references for low-quality trajectories, further improving the accuracy of their state estimation. This module increased MOTA to 64.2, IDF1 to 75.4, and reduced IDSW to 286, confirming the value of group behavior information in weak observation scenarios. Finally, when the GMCS fails, the STMP uses historical trajectory information to predict the future state of low-quality trajectories, effectively enhancing the temporal robustness of the model. This module further increased MOTA to 64.6, IDF1 to 76.1, and reduced IDSW to 276, indicating its ability to maintain target identity continuity in complex scenes.
Table 2: Complete results of the ablation study on the UAVDT dataset.